OpenStudy (anonymous):

hey im trying to find the solution set for the equation (*square root of x+1)+5=0. i keep arriving a -26 but apparently the answer is 0

OpenStudy (blockcolder):

No, the answer's not 0. There's no answer.

OpenStudy (anonymous):

its multiple choice and thats an option but my teacher says its 0 how did you arrive at no solution

OpenStudy (blockcolder):

First of all, if you plug in x=0, \(\sqrt{0+1}+5=6 \neq 0\), so it's wrong. I knew that there was no solution because: \[\sqrt{x+1}=-5\\ x+1=25\\ x=24\] but \(\sqrt{24+1}+5=10\), so x=24 also doesn't work. Since the possible solution did not satisfy the problem and we found no other possibility, that means that there are in fact no solutions.

OpenStudy (anonymous):

hm thats strange because like i said my teacher says its 0

OpenStudy (blockcolder):

Did he say why it's 0?

OpenStudy (anonymous):

no she was just reading off the answers i guess ill have to take the problem to her and ask her to explain

OpenStudy (blockcolder):

You should. :D

OpenStudy (anonymous):

thank you for the help how you worked the problem really made sense